Jump to main content or area navigation.

Contact Us

Waste Site Cleanup & Reuse in New England

Atlas Tack Corporation Superfund Site
Weekly Remedial Activities Update
Report # 19: November 13 - November 17, 2006

« Back to Atlas Tack Superfund Site Fact Sheet

1. ACTIVITIES PERFORMED DURING THE WEEK OF NOVEMBER 13 - NOVEMBER 17

  • Complete the Transport and disposal (T&D) of oversize material originating from stockpiles in the East Fill Area
  • Continued to excavate soils in the NW-Fill Area
  • Continued to screen soils originating from the NW-Fill Area
  • Consolidate stockpiles with same waste determination
  • Mobilize storage container and (2) 1,000-gal poly tanks from the New Bedford Superfund Site

Ongoing Site Activities:

  • Collected waste characterization soil samples from soils originating from NW-Fill Area for off site landfill profiling and disposal
  • Collected confirmatory soil samples from the floors of the NW-Fill Area
  • Continue water handling activities
  • Continued to inspect trucks prior to exiting the site
  • Continued to provide flaggers to cross trucks across the bike paths along Adams and Tripp Streets
  • Continued stockpile management activities
  • Continue with air monitoring program

Beginning November 30, 2006, biweekly meetings and updates will occur replacing the weekly meeting/updates. This change was made due to operations proceeding very well. Oversight coverage of operations will continue as they are now.

2. PROJECTED ACTIVITIES FOR THE WEEK OF NOVEMBER 20 - NOVEMBER 24

  • Observe Thanksgiving Holiday on 23rd and 24th of November.
  • Continue to T&D Oversize Debris (1,250-cy) to the Bourne Landfill, (500-cy) of COMM-97 soils to Marion, and (1,800-cy) of Sub-D soils to TREE (Turnkey NH landfill)
  • Continue with excavation activities in the Northwest Fill Area
  • Collect characterization soil samples from soils originating from Northwest Fill Area
  • Collect confirmatory soil samples from the NW-Fill Area
  • Screen soils originating from the Northwest Fill Area

Ongoing Site Activities:

  • Continue to collect pre and post excavation survey data of the excavations
  • Continue with stockpile management activities
  • Continue water handling activities
  • Continued to inspect trucks prior to exiting the site
  • Continued to provide flaggers to cross trucks across the bike paths along Adams and Tripp Streets
  • Continue with air monitoring program

3. AIR MONITORING INFORMATION

WESTON monitored for particulate (dust) during intrusive activities performed during this reporting period. WESTON recognizes EPA's level of concern for "PM10" to caution sensitive groups, including children, the elderly, and people with heart or lung disease, about exposure to coarse dust particles. The PM10 level of concern for sensitive groups is 150 µg/m3 for 24-hour average exposure.

Perimeter Air Monitoring is performed at up to (6) locations throughout a given work day. Readings are collected at 60 second intervals throughout the day and reported as a daily average.

If exceedances are identified, an increased effort of engineering controls (water suppression) will be utilized to adhere to the PM10 level of concern. If engineering controls appear to be inadequate, any dust generating activities will stop. Current dust control practices will be re-evaluated until concentrations are reduced to below the PM10 standard.

Results for 9 November – 15 November 2006 can be reviewed in the table presented below.

Weekly Dust Monitoring Summary
Week #19: Results for Thursday through Wednesday**

The EPA's PM10 level of concern for sensitive groups is 150 µg/m3 for 24-hour average exposure

Date Pleasant Street (µg/m3) Adams Street (µg/m3) Tripp Street (µg/m3) Egypt Road (µg/m3) Church Street (µg/m3)
11/9/2006 0 0 0 0 68
11/10/2006 0 0 0 0 0
11/13/2006 Precip Precip Precip Precip Precip
11/14/2006 Precip Precip Precip Precip Precip
11/15/2006          

Notes:
All readings reflect the daily average calculated over the course of the workday.
All readings collected with a Thermo Electron PDR 1000.
NSC: No sampling conducted; No intrusive work conducted in that area.
Precip: No sampling conducted due to precipitation.
*: Monitoring conducted over 24-hour period.
**: No work and air monitoring occurs on weekend days

4. T&D "WEEKLY" SUMMARY

DATE
DESTINATION
WASTE DETERMINATION
CY
TONS
STOCKPILES
17 Nov. 06 BOURNE Subtitle-D Oversize *319.93 *383.92 EA-15, EA-27, EA-21
16 Nov. 06 BOURNE Subtitle-D Oversize *319.93 *383.92 EA-15, EA-27, EA-21
15 Nov. 06 BOURNE Subtitle-D Oversize 319.33 383.20 EA-15, EA-27, EA-21
14 Nov. 06 BOURNE Subtitle-D Oversize 377.39 452.87 EA-15, EA-27, EA-21
13 Nov. 06 BOURNE Subtitle-D Oversize 354.10 424.92 EA-15, EA-27, EA-21
10 Nov. 06 BOURNE Subtitle-D Oversize 228.89 274.67 EA-15, EA-27, EA-21
* Projected based on daily average 11/10 – 11/15

5. T&D TOTAL SUMMARY

WK/E
DESTINATION
WASTE DETERMINATION
ORIGIN
CY
TONS
17 Nov. 06 Bourne Sub-D Oversize CID/ EAST FILL 1,919.58 2,303.5
10 Nov. 06 TREE SUB-D CID 1,130.37 1,695.56
03 Nov. 06 Marion COMM-97 CID/EAST FILL 1,881.51 2,822.27
27 Oct. 06 T&D DID NOT OCCUR THIS WEEK
20 Oct. 06 T&D DID NOT OCCUR THIS WEEK
13 Oct. 06 T&D DID NOT OCCUR THIS WEEK
06 Oct. 06 Marion COMM-97 CID 843.04 1,264.56
29 Sept. 06 Marion COMM-97 CID 449.63 674.45
22 Sept. 06 Crapo COMM-97 CID 291.23 436.85
15 Sept. 06 Crapo COMM-97 CID 533.67 800.51
T&D TOTAL
7,049.05 9,997.70

Jump to main content.